Circular No. 6451

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1996ao AND 1996ap IN ANONYMOUS GALAXIES

The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC 6418) reports the discovery of two apparent supernovae on CCD images taken by S. Chan on the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1996ao is located near Abell 3128 at α = 3^{h}28^{m}03^{s}.60, δ = −52^{o}53′53′′.6 (equinox 2000.0), and is offset 1′′.8 east and 0′′.8 south of the nucleus of the host galaxy. SN 1996ao is visible at mag ∼ 19 on V and R CCD images taken on Aug. 2, and was not detected on similar frames taken on June 24; a nearby star of mag 14 has position end figures 07^{s}.19, 56′08′′.7. SN 1996ap is located near Abell 3806 at α = 21^{h}45^{m}53^{s}.49, δ = −57^{o}08′27′′.3, and is offset 8′′.6 east and 4′′.0 north of the nucleus of the host galaxy. SN 1996ap is visible at mag ∼ 19.5 on V and R CCD images taken on Aug. 11, and was not detected on similar frames taken on Aug. 2; a nearby star of mag 13 has position end figures 36^{s}.63, 03′′.0. CCD Images taken with the MSSSO 2.3-m reflector on Aug. 14 show both supernovae at apparent brightnesses similar to those on the discovery frames. Spectra (range 350–1000 nm) obtained on Aug. 14 with the MSSSO 2.3-m Imager + Grism show the host galaxies’ redshifts for SN 1996ao and SN 1996ap, measured from Hα emission at the positions of the supernovae, to be z = 0.060 and z = 0.075, respectively. These are coincident with the redshifts of the host clusters, Abell 3128 (z = 0.060) and Abell 3806 (z = 0.076).

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude and coma-diameter estimates (cf. IAUC 6436):

July 21.03 UT, 5.5, 15′ (J.-C. Merlin, Le Creusot, France, 7×50 binoculars); Aug. 3.93, 5.6, 15′ (M. V. Zanotta, Mt. Ponna, Italy, 7×42 binoculars); 9.87, 5.3, 20′ (K. Hornoch, Lelekovice, Czech Republic, naked eye); 11.10, 5.5, 13′ (J. E. Bortle, Stormville, NY, 10×50 binoculars; tail 1^{o}.8 long in p.a. 130^{o}); 12.90, 5.5, 17′ (A. Pereira, Carnaxide, Portugal, 9×34 binoculars); 15.29, 5.2, 0^{o}.5 (S. J. O’Meara, Volcano, HI, naked eye).

4U 0115+634 Corrigendum. On IAUC 6450, line 13, for 1996 Aug. 5.5 UT.” read

1996 Aug. 5.583 TDB.”

1996 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green